If a ball weighs 115 N on planet Maricondia and has a mass of 60 kg on planet Earth. What is the acceleration due to the gravitational pull of planet Maricondia?



1
Expert's answer
2020-12-04T07:32:52-0500
"g_M=\frac{W_M}{m}\\g_M=\frac{115}{60}=1.92\frac{m}{s^2}"